OpenStudy (anonymous):

9+-3√33 --------- -18

OpenStudy (anonymous):

quadratic equations

OpenStudy (anonymous):

why can't you just add 3 to 9? why do you have to divide?

OpenStudy (anonymous):

is it because 3 is being multiplied by the square root of 33 or is it because 3 has a radical sign by it?

OpenStudy (anonymous):

btw this was the process of solving the quadratic equation. the original problem was

OpenStudy (anonymous):

-10x^2-9x+1

OpenStudy (anonymous):

You can't add the 3 because it's being multiplied by the square root of 33.

OpenStudy (anonymous):

so only because it's being multiplied?

OpenStudy (anonymous):

Wait is the equation meant to be 9 plus or minus the square root of 33?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

9 +- 3 square root of 33

OpenStudy (anonymous):

Oh yeah, my bad. When you have two things being multiplied and something added, you have to go by the order of operations. Multiplication comes before adding. So you multiply \[3 * \sqrt{33} = 17.2336\] Then you add 9 and subtract 9. Two different equations \[17.2336 - 9 = 8.2336\] \[17.2336 + 9 = 26.2336\] Divide both by -18. The answers should be -.457 and -1.457

OpenStudy (anonymous):

uhmm... without it being put into a decimal

OpenStudy (anonymous):

by just leaving the radicals under a square sign

OpenStudy (anonymous):

without finding their square root

OpenStudy (anonymous):

im guessing it's because it's 3 being multiplied by the square root of 33